Complex Numbers
 
1+z+z2+z3+....zn-1=(z-a1)(z-a2)(z-an-1)
 
HOW??

Grade:12

1 Answers

Jayant Kishore
8 years ago
Since 1, a1 , a2 , .......  are the roots of zn-1=0 therefore
zn-1 = (z-1)(z-a1)(z-a2)(z-a3).......(z - an-1)
=>  (zn-1)/(z-1) = (z-a1)(z-a2)(z-a3)....(z - an-1
=>  1+z+z2+z3+...... zn-1 =    (z-a1)(z-a2)(z-a3)....(z - an-1)    [ since (zn-1)/(z-1) is the sum of Geometric                                                                                                 Progression with 1st term 1 and common ratio z]
Hence Proved